Given πœ‡ = 45, and 𝜎 = 5.5. (7 points)

a. What is the raw score when 𝑧 = βˆ’1.57?


b. What is the raw score when 𝑧 = 2.09?


c. What is the raw score when βˆ’0.48 < 𝑧 < 1.4?


d. What is the raw score when βˆ’2.17 < 𝑧 < 1.79? e. What is the raw score when 𝑧 = 0.09?


1
Expert's answer
2022-05-04T12:04:12-0400

a.


"z =\dfrac{x-\mu}{\sigma}= −1.57"

"x=45-1.57(5.5)=36.365"

b.


"z =\dfrac{x-\mu}{\sigma}= 2.09"

"x=45+2.09(5.5)=56.495"

c.


"z =\dfrac{x-\mu}{\sigma}= −0.48"

"x=45-0.48(5.5)=42.36"




"z =\dfrac{x-\mu}{\sigma}= 1.4"


"x=45+1.4(5.5)=52.7"

"42.36<x<52.7"

d.


"z =\dfrac{x-\mu}{\sigma}= −2.17"

"x=45-2.17(5.5)=33.065"




"z =\dfrac{x-\mu}{\sigma}= 1.79"


"x=45+1.79(5.5)=54.845"

"33.065<x<54.845"

e.


"z =\dfrac{x-\mu}{\sigma}= 0.09"

"x=45+0.09(5.5)=45.495"